You own a company that competes with Old World DVD Company. Instead of selling DVDs, however, your company sells music downloads from a Web site. Things are going well now, but you know that it is only a matter of time before someone comes up with a better way to distribute music. Your company just paid a $1.97 per share dividend, and you expect to increase the dividend 10 percent next year. However, you then expect your dividend growth rate to begin going down—to 5 percent the following year, 2 percent the next year, and to -3 percent per year thereafter. Based upon these estimates, what is the value of a share of your company’s stock? Assume that the required rate of return is 13 percent. (Round dividends in intermediate calculations to 4 decimal places, e.g. 1.5325 and final answer to 2 decimal places, e.g. 15.25.)

Diaz Corp. is expected to grow rapidly at a rate of 35 percent for the next seven years. The company's first dividend, to be paid three years from now, will be $5. After seven years, the company (and the dividends it pays) will grow at a rate of 7.75 percent. What is the value of Diaz stock with a required rate of return of 14 percent? (Round intermediate calculations to 3 decimal places, e.g. 15.251 and final answer to 2 decimal places, e.g. 15.20.)

Tin-Tin Waste Management, Inc., is growing rapidly. Dividends are expected to grow at rates of 30 percent, 35 percent, 25 percent, and 18 percent over the next four years. Thereafter, management expects dividends to grow at a constant rate of 7 percent. The stock is currently selling at $47.30, and the required rate of return is 15.0 percent. Compute the dividend for the current year (D0). (Round intermediate calculations and final answer to 2 decimal places, e.g. 15.20.)

UNC Bancorp has issued preferred stock with no maturity date. It has a par value of $100 and pays a quarterly dividend of $2.25. If the required rate of return is 8 percent, what is the value of the stock today? (Round answer to 2 decimal places, e.g. 52.75.)
Answer #1

Solution:-

DPS0 = $1.97

DPS1 = DPS0 (1+G)

DPS1 = $1.97 * (1+0.10)

DPS1 = $2.167

DPS2 = DPS1 (1+G)

DPS2 = $2.167 * (1+0.05)

DPS2 = $2.275

DPS3 = DPS2 * (1+G)

DPS3 = $2.275 * (1+0.02)

DPS3 = $2.321

DPS4 = DPS3 * (1+G)

DPS4 = $2.321 * (1-0.03)

DPS4 = 2.251

Terminal Value at the end of Year 4-

Terminal Value = \frac{DPS4 * (1+G))}{Ke-G}

Terminal Value = \frac{2.251 * (1-0.03))}{0.13+0.03}

Terminal Value = $13.65

To Calculate Value of Share-

А. E 1 Value of Stock Year 2 3 4 5 1 2 3 4 Dividend 2.167 2.275 2.321 15.901 Required Rate @ 13% 0.885 0.783 0.693 0.613 Pres

A B D Value of Stock (Formula Sheet) 2 3 Year 4 1 5 2 63 7 4 8 9 10 Dividend 2.167 2.275 2.321 =2.251+13.65 Required Rate @ 1

Value of Share is $15

As per HOMEWORKLIB POLICY we need to answer only one question at once so please ask other as separate one.

If you have any query related to question then feel free to ask me in a comment.Thanks.